Good day,

For the reaction
X(g)+3Y(g)⇌3Z(g)
Kp = 1.24×10−5 at a temperature of 169 ∘C
Calculate the value of Kc

Thanks
.
1
Expert's answer
2020-08-06T03:55:06-0400

Kp = Kc×(R×T)Δn

Kc = Kp/(R×T)Δn

There are 3 mol of gas on the product side of the chemical equation, and there are 1+3=4 mol of gaseous reactants, so Δn = (3-4) = -1 in this case.

T = 169 + 273.15 = 442.15 K

R = 0.08206 L×atm/mol×K

Kc = 1.24×10-5/(0.08206×442.15)-1 = 1.24×10-5/(0.08206×442.15)-1 = 1.24×10-5/(36.28)-1 = 36.28×1.24×10-5 = 44.98×10-5 = 4.49×10-4

Answer: 4.49×10-4